# Promise and Peril: Stellar Contamination and Strict Limits on the Atmosphere Composition of TRAPPIST-1c from JWST NIRISS Transmission Spectra

### v1

@@ Abstract
Attempts to probe the atmospheres of rocky planets around M dwarfs present both promise and peril. While their favorable planet-to-star radius ratios enable searches for even thin secondary atmospheres, their high activity levels and high-energy outputs threaten atmosphere survival. Here, we present the 0.6–2.85 m transmission spectrum of the 1.1 , 340 K rocky planet TRAPPIST-1 c obtained over two JWST NIRISS/SOSS transit observations. Each of the two spectra displays 100–500 ppm signatures of stellar contamination. Despite being separated by 367 days, the retrieved spot and faculae properties are consistent between the two visits, resulting in nearly identical transmission spectra. Jointly retrieving for stellar contamination and a planetary atmosphere rules out with high confidence (3-) not only clear hydrogen-dominated atmospheres, but even thin, 1-bar high-mean molecular weight atmospheres rich in H2O, NH3, or CO (at the 2- level). We find that the only atmosphere scenarios which our spectrum cannot rule out are CH4- or CO2-rich atmospheres, which are both unlikely to be retained when considering the photodestruction of CH4 and the susceptibility of even a CO2-rich atmosphere to escape given the cumulative high-energy irradiation experienced by the planet. Our results further stress the importance of robustly accounting for stellar contamination when analyzing JWST observations of exo-Earths around M dwarfs, as well as the need for high-fidelity stellar models to search for the potential signals of thin secondary atmospheres.
@@ 1 Introduction
In recent years, the focus of such studies has been on small planets orbiting M dwarf stars, as the small size of these late-type host stars boosts the planet’s atmosphere signal, either in transmission or emission, compared to Sun-like stars (e.g., De Wit et al., 2016; Wakeford et al., 2019; Kreidberg et al., 2019; Lim et al., 2023; May et al., 2023; Greene et al., 2023; Zhang et al., 2024; Kirk et al., 2024). However, the potential promise of increased atmosphere observability also comes with a corresponding peril for atmosphere retention. Late-type stars have significantly longer pre-main sequence phases than Sun-like stars, during which time, orbiting planets are subjected to large amounts of high-energy radiation (Wordsworth & Pierrehumbert, 2013; Wheatley et al., 2017). Even at late times, during the M dwarf’s main sequence phase, they still maintain large outputs of high energy radiation, with ratios significantly higher than those of earlier-type stars (Pizzolato et al., 2003; Wright et al., 2011; Wheatley et al., 2017; Loyd et al., 2021), reaching up to 2000 that of the Sun for ultra-cool M dwarfs like TRAPPIST-1 (Wheatley et al., 2017). The increased amounts of high energy radiation received by M dwarf planets results in the potential for corresponding increases in the efficiency of atmosphere loss processes (e.g., Wordsworth & Pierrehumbert, 2013; Luger & Barnes, 2015; Bolmont et al., 2017; Turbet et al., 2023), resulting in uncertainty as to whether rocky planets around M dwarf stars can retain atmospheres to the present day (Zahnle & Catling, 2017).
@@
Moreover, late-type stars also maintain higher levels of activity and photospheric inhomogeneities (i.e., the presence of spots and faculae) than Sun-like stars (Peacock et al., 2019; Loyd et al., 2021). Unocculted spots and faculae on the host star’s photosphere (i.e., inhomogeneities that lie outside of a planet’s transit chord), in particular, are a pernicious problem preventing the detection of M-Earth atmospheres via transit observations (Rackham et al., 2018, 2019; TRAPPIST-1 JWST Community Initiative et al., 2024). The presence of unocculted inhomogeneties, also known as the transit light source effect (TLSE), can impart spurious features in the transmission spectra of rocky planets, which can often be significantly larger than expected atmosphere features (Rackham et al., 2018, 2023; Lim et al., 2023; TRAPPIST-1 JWST Community Initiative et al., 2024). Despite these challenges, transit observations using the Hubble Space Telescope (HST) and now with JWST, have been able to widely rule out cloud-free, H/He-dominated atmospheres for rocky M dwarf planets (De Wit et al., 2016; Wakeford et al., 2019; Libby-Roberts et al., 2022; Lim et al., 2023; Lustig-Yaeger et al., 2023; Cadieux et al., 2024; Damiano et al., 2024; Kirk et al., 2024), although such extended atmospheres are already disfavoured based on mass and radius measurements alone (Agol et al., 2021). In many cases though, the TLSE, still prevents us from robustly probing secondary, or high-mean molecular weight atmospheres in transmission (e.g., Wakeford et al., 2019; Lim et al., 2023; Cadieux et al., 2024; May et al., 2023; Moran et al., 2023).
@@
Observations of M-dwarf planets in eclipse geometry circumvent the TLSE problem but bring with them their own sets of challenges: the planet’s orbital eccentricity must be well known to calculate the phase of the eclipse, emission observations favour planets with high temperatures, potentially ruling out habitable-zone worlds, etc. Though, even without the impacts of the TLSE, emission studies of rocky M-dwarf planets, both in the JWST era and previously, have either been negative (Kreidberg et al., 2019; Greene et al., 2023; Ih et al., 2023; Zhang et al., 2024; Mansfield et al., 2024; Xue et al., 2024, i.e., favour atmosphereless interpretations;) or unable to provide definitive evidence for the presence of a high-mean molecular weight atmosphere (e.g., Zieba et al., 2023).
@@
Here, we present transit observations of TRAPPIST-1 c with JWST NIRISS/SOSS. TRAPPIST-1 c is the second innermost planet (=0.01580 AU, =340 K) of the TRAPPIST-1 system, with a mass and radius of 1.3 and 1.1 that of the Earth, respectively (Gillon et al., 2016; Ducrot et al., 2020; Agol et al., 2021). It was first observed in transit by De Wit et al. (2016) with HST/WFC3 from 1.1 – 1.7 µm. Their observations rule out cloud-free, H/He-dominated atmosphere scenarios at 10-, but lack the precision to constrain the presence of a wide range (e.g., H2O-rich, Venus-like) of higher-mean molecular weight atmospheres. TRAPPIST-1 c was then observed in emission with JWST MIRI/F1500W photometry by Zieba et al. (2023) at 15 µm. Their dayside brightness temperature of 38031 K rules out thick (10 bar), CO2-rich atmospheres, but remains consistent with a thin (0.01 bar) atmosphere composed purely of CO2, or trace amounts of CO2 in thinner (0.1 – 10 bar), O2-dominated atmospheres at a 2 level.
@@
This paper is organized as follows. We present the observations and data analysis in Section 2. We describe our analysis of the TLSE in Section 3, and joint modelling of the atmosphere and stellar contamination in Section 4. We then discuss our results in Section 5 and follow up with a short conclusion in Section 6.
@@ 2.1 Data Reduction
We reduce the time series observations (TSO) with the exoTEDRF pipeline111https://github.com/radicamc/exoTEDRF, formerly known as supreme-SPOON. (Radica, 2024; Feinstein et al., 2023; Radica et al., 2023), closely following standard procedures (e.g., Radica et al., 2024; Cadieux et al., 2024). We perform the correction of 1/ noise at the group-level (that is, before ramp fitting) using the scale-achromatic-window method introduced in exoTEDRF v1.4.0. With this algorithm, we only use rows within 30 pixels of the trace to estimate the 1/ level, which can potentially help to reduce residual red noise in SOSS TSOs, particularly with high group numbers (e.g., Feinstein et al., 2023; Holmberg & Madhusudhan, 2023). We also perform a “piecewise” background subtraction (e.g., Lim et al., 2023; Fournier-Tondreau et al., 2024; Radica et al., 2024), whereby the standard STScI SOSS background model is scaled separately on either side of the background “step”, as we found that a single scaling for the whole detector did not adequately remove the background signal.
@@ 2.2 White Light Curve Fitting
We construct white light curves for each visit and each order by summing all flux on the detector in order 1 (0.85 – 2.8 µm), and only wavelengths 0.85 µm in order 2. We fit these white light curves using the flexible juliet package (Espinoza et al., 2019). For each visit, we jointly fit the first and second order white light curves, as in Radica et al. (2024). Concretely, we fit an astrophysical transit model, as computed by BATMAN (Kreidberg, 2015), with the orbital parameters (, the time of mid-transit, , the orbit impact parameter, and , the scaled orbital semi-major axis) shared between both orders, and the scaled planet radius, fit to each order individually. We, furthermore, assume a circular orbit and fix the orbital period to 2.421937 d (Agol et al., 2021). We fit, separately for each order, the two parameters of the quadratic limb darkening law, following the Kipping (2013) parameterization. We also fit a systematics model to the light curve for each order. For visit 2, the systematics model is composed of a linear trend with time, whereas for visit 1 we also include a quadratic term with time. TRAPPIST-1 is well-known to be a highly active star (Lim et al., 2023), and we, therefore, include a Gaussian process (GP) with a simple harmonic oscillator (SHO) kernel to model residual red noise in each light curve not captured by the aforementioned detrending. The SHO kernel, as implemented in celerite (Foreman-Mackey et al., 2017) has proven efficient at modelling the impacts of host star photospheric variations on light curves (i.e., stellar granulation, Kallinger et al., 2014; Pereira et al., 2019; Radica et al., 2024). For each visit, we fit the characteristic amplitude of the GP separately to each order but share the characteristic frequency between the two (e.g., Radica et al., 2024). Finally, we include an error inflation term for each order. Our final models therefore consist of 20 parameters for visit 1 and 18 for visit 2 — we use wide and uninformative priors in all cases.
@@
The results of the white light curve fits are shown in Figure 1. During the pre-transit baseline of the second visit, there was a large flare event on the host star. We attempted several methods to correct this flare in the light curves e.g., detrending against the H flux, computed for each integration by measuring the area between the flux in the H line and a local quadratic continuum, shown in the second panels in Figure 1. However, this avenue was unsuccessful as there is a slight time delay between the appearance of the flare in the H time series and in the white light curve, likely due to lags in energy release as a function of wavelength (Howard et al., 2023). We eventually decided that, since the flare occurred in the pre-transit baseline and was sufficiently separated in time from the transit itself, the optimal course of action was to simply cut the affected integration from the fit. For visit 2, we, therefore, cut the first 75 integrations of the TSO (denoted with grey shading in Figure 1).
@@ 2.3 Spectroscopic Light Curve Fitting
To construct the spectrophotometric light curves, we sum the first order flux in bins of 80 pixels (corresponding to a constant resolving power of approximately =25). The combined effects of the spectrum of the late-M host star, and the low throughput of the SOSS order 2 mean that the second order trace is incredibly faint for observations of TRAPPIST-1 planets (e.g., Lim et al., 2023). We therefore do not bin order 2 further, and simply include the white light curve results in our transit spectrum. The orbital parameters from both visits are consistent between visits (Table 1), and we therefore fix them to the weighted average between the two when fitting the spectroscopically binned light curves. We include the same systematics model as mentioned above for each visit, except that we fix the characteristic timescale of the GP to the best-fitting value from the white light fit for each visit, and allow the amplitude to freely vary, since we can reasonably expect the impacts of stellar variability to be more prominent at bluer vs. redder wavelengths (e.g., Radica et al., 2024).
We find a slight offset of 122 ppm between the spectra of of the two visits, which we attribute to residual dilution from the pre-transit flare in visit 2. Indeed, this can also be noticed in the white light curve depths for each visit shown in Table 1, where the visit 2 order 1 transit depth is 100 ppm shallower than visit 1 — suggestive of dilution. We tested trimming a different number of pre-transit integrations and found that this slightly changed the offset (to the tune of 5 – 10 ppm), but did not change the shape of the final transit spectrum.
@@
We also reduce and fit the TSOs from each visit using the independent NAMELESS pipeline (Feinstein et al., 2023; Coulombe et al., 2023). Further details of this analysis are presented in Appendix A, and a comparison of the spectra from both pipelines are shown in Figure 6. Since the spectra from each pipeline are consistent to within 1- for both visits, we only consider the exoTEDRF spectrum for the remainder of our analysis.
3 Modelling of the transit light source effect

### v3

@@ Abstract
Attempts to probe the atmospheres of rocky planets around M dwarfs present both promise and peril. While their favorable planet-to-star radius ratios enable searches for even thin secondary atmospheres, their high activity levels and high-energy outputs threaten atmosphere survival. Here, we present the 0.6–2.85 m transmission spectrum of the 1.1 , 340 K rocky planet TRAPPIST-1 c obtained over two JWST NIRISS/SOSS transit observations. Each of the two spectra displays 100–500 ppm signatures of stellar contamination. Despite being separated by 367 days, the retrieved spot and faculae properties are consistent between the two visits, resulting in nearly identical transmission spectra. Jointly retrieving for stellar contamination and a planetary atmosphere reveals that our spectrum can rule out hydrogen-dominated, 300 solar metallicity atmospheres with effective surface pressures down to 10 mbar at the 3- level. For high-mean molecular weight atmospheres, where O2 or N2 is the background gas, our spectrum disfavors partial pressures of more than 10 mbar for H2O, CO, NH3 and CH4 at the 2- level. Similarly, under the assumption of a 100% H2O, NH3, CO, or CH4 atmosphere, our spectrum disfavors thick, 1 bar atmospheres at the 2- level. These non-detections of spectral features are in line with predictions that even heavier, CO2-rich, atmospheres would be efficiently lost on TRAPPIST-1 c given the cumulative high-energy irradiation experienced by the planet. Our results further stress the importance of robustly accounting for stellar contamination when analyzing JWST observations of exo-Earths around M dwarfs, as well as the need for high-fidelity stellar models to search for the potential signals of thin secondary atmospheres.
@@ 1 Introduction
In recent years, the focus of such studies has been on small planets orbiting M dwarf stars, as the small size of late-type stars boosts the planet’s atmosphere signal, either in transmission or emission, compared to Sun-like stars (e.g., De Wit et al., 2016; Wakeford et al., 2019; Kreidberg et al., 2019; Lim et al., 2023; May et al., 2023; Greene et al., 2023; Zhang et al., 2024; Kirk et al., 2024).
@@
However, the potential promise of increased atmosphere observability also comes with a corresponding peril for atmosphere retention. Late-type stars have significantly longer pre-main sequence phases than Sun-like stars, during which time orbiting planets are subjected to large amounts of high-energy radiation (Wordsworth & Pierrehumbert, 2013; Wheatley et al., 2017). Even at late times, during the M dwarf’s main sequence phase, they still maintain large outputs of high energy radiation, with ratios significantly higher than those of earlier-type stars (Pizzolato et al., 2003; Wright et al., 2011; Wheatley et al., 2017; Loyd et al., 2021), reaching up to 2000 that of the Sun for ultra-cool M dwarfs like TRAPPIST-1 (Wheatley et al., 2017; Bourrier et al., 2017). The increased amounts of high energy radiation received by M dwarf planets results in the potential for corresponding increases in the efficiency of atmosphere loss processes (e.g., Wordsworth & Pierrehumbert, 2013; Luger & Barnes, 2015; Bolmont et al., 2017; Turbet et al., 2023), resulting in uncertainty as to whether rocky planets around M dwarf stars can retain atmospheres to the present day (Zahnle & Catling, 2017).
@@
Moreover, late-type stars also maintain higher levels of activity and photospheric inhomogeneities (i.e., the presence of spots and faculae) than Sun-like stars (Peacock et al., 2019; Loyd et al., 2021). Unocculted spots and faculae on the host star’s photosphere (i.e., inhomogeneities that lie outside of a planet’s transit chord), in particular, are a pernicious problem preventing the detection of atmospheres around Earth-like planets orbiting M dwarfs via transit observations (Rackham et al., 2018, 2019; Wakeford et al., 2019; Garcia et al., 2022; TRAPPIST-1 JWST Community Initiative et al., 2024; Rackham & de Wit, 2024). The presence of unocculted inhomogeneities can impart spurious features on the transmission spectra of rocky planets, which can often be significantly larger than expected atmosphere features (Rackham et al., 2018, 2023; Lim et al., 2023; TRAPPIST-1 JWST Community Initiative et al., 2024; Rackham & de Wit, 2024). This phenomenon is often referred to as the transit light source effect (TLSE). Despite these challenges, transit observations using the Hubble Space Telescope (HST) and now with JWST, have been able to widely rule out cloud-free, H/He-dominated atmospheres for rocky M dwarf planets (De Wit et al., 2016; Wakeford et al., 2019; Garcia et al., 2022; Libby-Roberts et al., 2022; Lim et al., 2023; Lustig-Yaeger et al., 2023; Cadieux et al., 2024; Damiano et al., 2024; Kirk et al., 2024), although such extended atmospheres are often already disfavoured based on mass and radius measurements alone (e.g., Agol et al., 2021). In many cases though, the TLSE, still prevents us from robustly probing secondary, or high-mean molecular weight atmospheres in transmission (e.g., Wakeford et al., 2019; Lim et al., 2023; Cadieux et al., 2024; May et al., 2023; Moran et al., 2023).
@@
Observations of M dwarf planets in eclipse geometry circumvent the TLSE problem but bring with them their own sets of challenges: the planet’s orbital eccentricity must be well known to calculate the phase of the eclipse, emission observations favour planets with high temperatures potentially ruling out habitable-zone worlds, etc. Though, even without the impacts of the TLSE, emission studies of rocky M-dwarf planets, both in the JWST era and previously, have either been negative (Kreidberg et al., 2019; Greene et al., 2023; Ih et al., 2023; Zhang et al., 2024; Mansfield et al., 2024; Xue et al., 2024, i.e., favour atmosphereless interpretations;) or unable to provide definitive evidence for the presence of a high-mean molecular weight (MMW) atmosphere (e.g., Zieba et al., 2023; Lincowski et al., 2023), often due to degeneracies between atmospheric absorption and surface reflectivity (e.g., Ih et al., 2023; Lincowski et al., 2023).
@@
Here, we present transit observations of TRAPPIST-1 c with JWST NIRISS/SOSS. TRAPPIST-1 c is the second innermost planet (=0.01580 AU, =340 K) of the TRAPPIST-1 system, with a mass and radius of 1.3 and 1.1 that of the Earth, respectively (Gillon et al., 2016; Ducrot et al., 2020; Agol et al., 2021). It was first observed in transit by De Wit et al. (2016) with HST/WFC3 from 1.1 – 1.7 µm. Their observations rule out cloud-free, H/He-dominated atmosphere scenarios at 10-, but lack the precision to constrain the presence of a wide range (e.g., H2O-rich, Venus-like) of higher-MMW atmospheres.
TRAPPIST-1 c was then observed in emission with JWST MIRI/F1500W photometry by Zieba et al. (2023) at 15 µm. Their dayside brightness temperature of 38031 K rules out thick (10 bar), CO2-rich atmospheres, but remains consistent with a thin (0.01 bar) atmosphere composed purely of CO2, or trace amounts of CO2 in thinner (0.1 – 10 bar), O2-dominated atmospheres at a 2 level. This was then followed up by Lincowski et al. (2023) who compared the Zieba et al. (2023) eclipse depth to a broader range of potential atmosphere scenarios using a self-consistent modelling framework. Their results generally agree with those of Zieba et al. (2023), in that trace amounts of CO2 in a thin O2-dominated atmosphere provide the best matches to the data. However, they also also find that 0.1 bar pure-O2, and even 3 bar steam atmospheres are consistent with the data within 2. Venus-like atmospheres, though, are disfavoured at 3.
@@
This paper is organized as follows. We present the observations and data analysis in Section 2. We describe our analysis of the stellar photosphere in Section 3, and joint modelling of the atmosphere and stellar contamination in Section 4. We then discuss our results in Section 5 and follow up with a short conclusion in Section 6.
@@ 2.1 Data Reduction
We reduce the time series observations (TSO) with the exoTEDRF pipeline111https://github.com/radicamc/exoTEDRF (Radica, 2024; Feinstein et al., 2023; Radica et al., 2023), closely following standard procedures as described in, e.g., Radica et al. (2024); Cadieux et al. (2024); Piaulet-Ghorayeb et al. (2024). Concretely, we perform the standard exoTEDRF stage 1 calibrations, including the correction of 1/ noise at the group-level (that is, before ramp fitting) using the scale-achromatic-window method introduced in exoTEDRF v1.4.0. With this algorithm, we only use rows within 30 pixels of the trace to estimate the 1/ level, which can potentially help to reduce residual red noise in SOSS TSOs, particularly with high group numbers (e.g., Feinstein et al., 2023; Holmberg & Madhusudhan, 2023). We then proceed with the standard Stage 2 calibrations (e.g., flat field correction, bad pixel interpolation, etc.). Notably, we perform a “piecewise” background subtraction (e.g., Lim et al., 2023; Fournier-Tondreau et al., 2024; Radica et al., 2024), whereby the standard STScI SOSS background model is scaled separately on either side of the background “step”, as we found that a single scaling for the whole detector did not adequately remove the background signal. As a final step in the calibrations, we performed a principle component analysis (PCA) on the 2D detector images (e.g., Coulombe et al., 2023; Radica et al., 2024) to identify any detector-level trends (i.e., mirror tilt events, trace position drifts), which could impact the light curves.
@@ 2.2 White Light Curve Fitting
We construct white light curves for each visit and each order by summing all flux on the detector in order 1 (0.85 – 2.8 µm), and only wavelengths 0.85 µm in order 2. We fit these white light curves using the flexible exoUPRF library (Radica, 2024). For each visit, we jointly fit the first and second order white light curves, as in Radica et al. (2024). Concretely, we fit an astrophysical transit model, as computed by BATMAN (Kreidberg, 2015), with the orbital parameters (, the time of mid-transit, , the orbital inclination, and , the scaled orbital semi-major axis) shared between both orders, and the scaled planet radius, fit to each order individually. We, furthermore, assume a circular orbit and fix the orbital period to 2.421937 d (Agol et al., 2021). We fit, separately for each order, the two parameters of the quadratic limb darkening law, , in order to avoid potential biases that can be introduced by employing the Kipping (2013), parameterization (e.g., Coulombe et al., 2024). We also tested several other limb-darkening treatments but found that the choice of parameterization did not impact the final spectra.
We also fit a systematics model to the light curve for each order. For each visit, the first component of the systematics model is a linear trend with time. TRAPPIST-1 is well-known to be a highly active star (Lim et al., 2023), and we, therefore, include a Gaussian process (GP) with a Matérn 3/2 kernel, to model residual red noise in the visit 1 light curves not captured by the aforementioned detrending. We train the GP on the timeseries of the second eigenvalue extracted from the PCA, which roughly corresponds to sub-pixel drifts of the trace position on the detector, as we find this reduces the residual scatter in the light curves by a further 10% compared to training the GP on the light curve time axis. The characteristic amplitude of the GP was fit independently for each order, but the timescale was shared between both (e.g., Radica et al., 2024).
@@
During the pre-transit baseline of the second visit, there was a large flare event on the host star. We to correct this flare in the light curves by detrending against the H flux, computed for each integration by measuring the area between the flux in the H line and a local quadratic continuum, shown in the second panels in Figure 1. However, this avenue was unsuccessful as there is a slight time delay between the appearance of the flare in the H time series and in the white light curve, likely due to lags in energy release as a function of wavelength (Howard et al., 2023). However, given the broad structure of this flare compared to others seen in TRAPPIST-1 light curves (e.g., Howard et al., 2023), we found that we were able to adequately include the flare in our light curve fits using a GP which again employed the Matérn 3/2 kernel. As above, we fit the characteristic amplitude separately to each order, but shared the timescale between both.
Finally, we include an error inflation term for each order of each visit, which is added in quadrature to the extracted flux errors. Our final models consist of 20 parameters for each visit — we use wide and uninformative priors in all cases. The results of the white light curve fits are shown in Figure 1 and the best-fitting astrophysical parameters are summarized in Table 1.
@@ 2.3 Spectroscopic Light Curve Fitting
To construct the spectrophotometric light curves, we sum the first order flux in bins of 80 pixels (corresponding to a constant resolving power of approximately =25). The combined effects of the spectrum of the late-M host star, and the low throughput of the SOSS order 2 mean that the second order trace is very faint for observations of TRAPPIST-1 planets (e.g., Lim et al., 2023). We therefore do not bin order 2 further, and simply include the white light curve results in our transit spectrum. The orbital parameters from both visits are consistent between visits (Table 1) as well as with literature values (e.g., Agol et al., 2021), and so we fix them to the weighted average between the two visits when fitting the spectroscopically binned light curves. We include the same systematics model as mentioned above for each visit, except that we fix the characteristic timescale of the GP to the best-fitting value from the white light fit for each visit, and allow the amplitude to freely vary, since we can reasonably expect the impacts of stellar variability to be more prominent at bluer vs. redder wavelengths (e.g., Radica et al., 2024).
@@
We also reduce and fit the TSOs from each visit using the independent NAMELESS pipeline (Feinstein et al., 2023; Coulombe et al., 2023). Further details of this analysis are presented in Appendix A, and a comparison of the spectra from both pipelines are shown in Figure 7. Since the spectra from each pipeline are consistent to within 1- for both visits (or 1 via a -per-data-point metric), we only consider the exoTEDRF spectrum for the remainder of our analysis.
3 Modelling of the Stellar Photosphere
Before undertaking any planet atmosphere analysis, we first seek to better constrain the properties of TRAPPIST-1’s photosphere, and thereby aid in our attempts to disentangle veritable atmosphere signals from the TLSE. To this end, we perform two independent, but complementary analyses: one of the in-transit data via the TLSE (Section 3.1), and another focusing on the out-of-transit data via modelling the stellar spectrum itself (Section 3.2).
3.1 Transit Light Source Effect Modelling
. Stellar contamination alone can explain the wavelength-dependent variations in the spectrum, and the retrieved heterogeneity parameters are largely consistent across both visits.